OpenStudy (anonymous):

What is the exact value of the expression 7√8 + 2√12 - 3√2?

OpenStudy (anonymous):

Show your work.

OpenStudy (anonymous):

\[3\sqrt{2}-3\sqrt{2}+2\sqrt{2}\]

OpenStudy (anonymous):

2√2

OpenStudy (anonymous):

opss, bad conts :)

OpenStudy (anonymous):

7√8 = 14√2 2√12 = 4√3 3√2 11√2-4√3

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

sure do! :) thanks!